OpenStudy (anonymous):

Please Help... f(x)=2/x; g(x)=2/x Find f(g(x)) Find g(f(x)) x≠ ? I know it = 2 for both, but how do I write it?

OpenStudy (anonymous):

f(g(x))=2/(2/x)=x

OpenStudy (anonymous):

g(f(x))=2/(2/x)=x

OpenStudy (anonymous):

and x is not equal to 0

OpenStudy (anonymous):

Do I type it ast x≠0?

OpenStudy (anonymous):

yes, its the only possible correct answer